Please show all your work for the problems!

Exp

Composition

Total Vol

pH

1

A few mL of 0.10M NH3 (unchanged)

n/a

11.13

2

2.5mL 0.1M NH3 & 7.5mL H2O

10 ml

10.82

3

2.5mL 0.1M NH4Cl & 7.5mL H2O

10 ml

6.01

4

2.5mL 0.1M NH3 & 2.5mL 0.1M NH4Cl & 5mL H2O

10 ml

9.26

5

Take 1mL Solution #4 and add 9mL of water

10 ml

9.26

Solution 1. Use the concentration of NH3 and the measured pH above to determine the Kb of NH3. (show the ICE table and calculations)

Use the equation (Ka*Kb = 1*10-14) to calculate the Ka for NH4+ ?

Solution 2. Use the equation M1V1 = M2V2 to calculate new conc. of NH3.

Conc. of NH3 ? pH measured:    10.82

Compare solutions 1 & 2. What effect did diluting the base have on the pH?

Use the new concentration and pH to measure the Kb of NH3 again. Did it change?

Solution 3. Conc. of NH4+ ?    pH of solution 3: 6.01

Use the diluted concentration of NH4+ and measured pH to determine the Ka of NH4+

How does this Ka value compare to the value on pg. 1? (within a power of 10 is good)
Answer #1

Solution 1. pOH = 1/2pKb - 1/2 log(c)

So, 14-11.13 = 1/2pKb - 1/2 log(0.1). Hence, pKb = 4.74

Thus, Kb of NH3 = 1.8 x 10-5

pKa of NH4+ = 14 - 4.74 = 9.26

Thus, Ka of NH4+ = 5.4954 x 10-10

Solution 2. Concentration of NH3 = (2.5*0.1)/10 = 0.025 M

On Dilution of the Base, the [OH-] dropped, and thus the pH also dropped.

Again, 14-10.82 = 1/2pKb - 1/2 log(0.025)

So, pKb = 4.76. Thus, the pKb doesn't change significantly.

Kb = 1.746 x 10-5

​​​​

Solution 3. Concentration of NH4+ = (2.5*0.1)/10 = 0.025 M

Now, 6.01 = 1/2pKa - 1/2 log(0.025)

Thus, pKa = 10.42

Hence, Ka = 3.82 x 10-11

Thus, the value compares good according to the Solution 1.